OpenStudy (anonymous):

A.What is the fourth term of (3/2y^2+2/3y^-3)^7? B. Consider the binomial product (a+b)^n where the third term of the expansion is 127575x^5y^4 and the fourth term of the expansion is -354375x^4y^6. Find a,b, and n.

OpenStudy (campbell_st):

just to make life easy is the binomial \[(\frac{3}{2} y^2 + \frac{2}{3} y^{-3})^7\]

OpenStudy (anonymous):

Yes that is binomial (:

OpenStudy (campbell_st):

ok... so the 4th term is \[^7C_{3} (\frac{3}{2}y^2)^{7 -3} \times (\frac{2}{3} y^{-3})^3\] which becomes \[35 \times \frac{3^4}{2^4}(y^2)^4 \times \frac{2^3}{3^3}(y^{-3})^3\] you need to simplify it... as far as the fractions go you'll have \[\frac{3^4}{2^4} \times \frac{2^3}{3^3} = \frac{3}{2}\] hope it helps

OpenStudy (anonymous):

Thank you! This does help. How did you get the 35? and is the y with the negative exponent to a power of 3 or two?

OpenStudy (campbell_st):

well 35 is the 7C3 combination..... you can also find the coefficient of 35 from pascals triangle

OpenStudy (anonymous):

perfect. thank you!